One root of the equation 2x^3 – 5x^2 – 8x + d = 0 is the negative of another. Compute d.
(A) –10 (B) –4 (C) 8 (D) 20 (E) 24

Answer by Jk22(389)

(D) d=20
2(x^2-a^2)(x+b)=2x^3+2bx^2-2a^2x-2a^2b
b=-5/2, a=2, d=-2a^2b=20